OpenStudy (anonymous):

Graph the function. y = 7(6)x+2 + 1

OpenStudy (anonymous):

\[y=7(6)^{x+2} +1\]

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (dan815):

graph some easy points when x=-2 the exponent goes to 0 so you get y=7*6^0 + 1 =7*1+1 = 8

OpenStudy (dan815):